how many miles per hour would i need to run in order to finish 1 & 1/3 mile distance in 12 minutes
Answer by Alan3354(69443)

how many miles per hour would i need to run in order to finish 1 & 1/3 mile distance in 12 minutes
--------------
d = r*t
r = d/t = (4/3) miles/(1/5) hours